OpenStudy (anonymous):

Calculate the cross product assuming that u*w = <7,6,-4> (-5u + 3w)*w =<_______>

OpenStudy (helder_edwin):

do u mean \[\large u\times w=(7,6,-4) \] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ganeshie8 i need help

ganeshie8 (ganeshie8):

hint : cross product is distributive

ganeshie8 (ganeshie8):

\((-5u + 3w) \times w = -5u \times w + 3w \times w\)

OpenStudy (anonymous):

how do i substitute the values ?

ganeshie8 (ganeshie8):

\((-5u + 3w) \times w = -5u \times w + 3w \times w\) \( = -5 u \times w + 3 * 0\) \( = -5 u \times w \)

ganeshie8 (ganeshie8):

cross product of same vectors is 0 \(i \times i = 0\)

ganeshie8 (ganeshie8):

substitute the given \(u \times w\) value now

OpenStudy (anonymous):

=-5<7, 6 -4> =<-35, -30, 20>

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thanks alot

ganeshie8 (ganeshie8):

np :)